Sirs, I am seeking help with my sons Algebra as I am really struggling to comprehend the basics.
(Multiply by the denominator to eliminate the denominator, but remember to do so on both sides. Then eliminate the number in front of the letter by division.)
2x=4
3
The 1st numbers are fractions.
Very grateful for any help.

Answer by checkley75(3666)

2x/3=4 cross multiply
2x=4*3
2x=12
x=12/2
x=6 answer.